OpenStudy (anonymous):

If a baseball player hits a baseball from 4 feet off the ground with an initial velocity of 64 feet per second, how long will it take the baseball to hit the ground? Use the equation h = –16t2 + 64t + 4.

OpenStudy (ash2326):

h=-16t^2+64t+4 t=0 the ball is at height 4 feet this can also be verified from the given equation we have to find value of t for which h=0 0=-16t^2+64t+4 or 4t-16t-1=0 we know that ax^2+bx+c=0 \[x=(-b \pm \sqrt(b^2-4ac))/2a\] here a=4 b=-16 c=-1 so \[t=(16 \pm \sqrt(16^2-4*4*(-1))/8\] \[t=(16 \pm \sqrt(272))/8\] or \[t=(16 \pm 16 \sqrt 17)/8\] t= 2 + 2 root 17 or t= 2- 2 root 17 t= 2-2 root 17 <0 so our solution is t= 2+ 2 root 17 \[t=2(1+ \sqrt 17)\] t=10.2462 seconds

OpenStudy (anonymous):

t=(16±(√272))/8 t=(16±(√16*√17))/8 t=(16±(4√17))/8 and that's where I'm stuck